Can anyone provide me with how the author came up with the numbers for the "year one portfolio"? Really confused on how he got the 9600 in parentheses....Buying Your First Property Using these standards, we find a fourplex that needs a little help. Perhaps its a bank repo that smells like a hundred cats, or a for-sale-by-owner (FSBO) property owned by someone who cant quite take care of it. Im not talking about a disaster-just a place that needs some cosmetic assistance. The property is listed at $100,000, but we are able to buy it for $80,000. For simplicitys sake, well say that the seller will pay our closing costs, which means we are required to put down $16,000 for a 20% down payment. In addition, we will spend $4,000 on minor repairs and improvements to satisfy the standard of 10% appreciation during the first year Year One Here is a summary of our portfolio after year one: Our loan is for $63,500 (approximately). Our value is at $110,000 (10% appreciation during year one). Our cash flow saved (from cash flow) is $10,000. (OK, technically $9,600.I like easy math.) Our total equity is $46,500. Our total net worth is $56,500. .
Answer #1

1. =80000-16000-500(closing cost)=63500

2. =100000*(1+10%)=110000

3.

4. =110000(from 2 above)-63500(from 1 above)=46500

5. =46500+10000(from 3 above)=56500